OpenStudy (anonymous):

Find the length of the radius of a sphere with a surface area of 314.16 ft2. A. 5 ft B. 15 ft C. 10 ft D. 2.5 ft

OpenStudy (experimentx):

10

OpenStudy (anonymous):

10

OpenStudy (king):

5

OpenStudy (king):

its actually 4.998999899998...

OpenStudy (anonymous):

I concur with mr. King here. 4pi r^2 = 314.16 4r^2 = 100 r^2 = 25 r = 5